f(x)=x³+3x²-x-3; f(1) = 0

Answer by ikleyn(52905)

.

Factorize by grouping


    f(x) = x^3 + 3x^2 - x - 3 = (x^3 - x) + (3x^2 - 3) = x*(x^2-1) + 3*(x^2-1) = 


         = (x+3)*(x^2-1) = (x+3)*(x+1)*(x-1).



Since you didn't ask question, I could not answer.


But I prepared EVERYTHING for you in order for you be able to create your own answer.

--------------

Do not submit posts with no questions.

--------------

If your native language is not English and if it is difficult for you to formulate your question in English,
post it in your native language.

For leading world languages, we have Google-translator.